If it takes 4 seconds for hydrogen to diffuse from 1 and of a gas jar to another, how long will it take oxygen to diffuse through the same distance
1
Expert's answer
2020-06-22T09:15:42-0400

Rate of effusion "=" "Amount of gas \over time"

Combine with Grahams law;

"Rate of effusion of hydrogen\over Rate of effusion of Oxygen" "=" "\sqrt MO \over \sqrt MH"

To get

"{Amount of H transferred \over time of H} \over{Amount of O transfared\over time of O}" "=" "\sqrt MO \over \sqrt MH"

"\therefore" "Time of H\over Time of 0" "=" "\sqrt MO\over \sqrt MH" "=" "\sqrt MO\over \sqrt MH"


"Time of O\over 4s" "=" "\sqrt 32g/mol \over 2mol/L" "=" "\sqrt 16 \over \sqrt 1"

"Time of O \over 4s" "=" "4\over 1"

"=4*4=16"

Oxygen takes "16" seconds to travel through the same distance.